# Search for Dark Matter Annihilation to gamma-rays from SPT-SZ selected Galaxy Clusters

### v1

@@ IV.1 Data selection
We also applied the filters, DATA_QUAL > 0 and LAT_CONFIG == 1, which were utilized to eliminate low-quality data. The rock angle was constrained to abs(rock angle) and |b| to mitigate particle contamination. To reduce atmospheric interference, a zenith angle cut of 90∘ was imposed on the events. We also removed all clusters within of the galactic plane. The spatial binning of the data was performed at a resolution of 0.2 pixels, which helps to reduce noise and improve the signal-to-noise ratio. Our background model incorporated all the gamma-ray sources from the fourth Fermi-LAT catalogue (4FGL-DR4) Ballet et al. (2023), encompassing both point-like and extended sources. To account for the pervasive, low-intensity gamma-ray emission from the Milky Way galaxy itself (diffuse emission), we employed the Galactic diffuse emission model (gll_iem_v07.fits). Additionally, an isotropic component (iso_P8R3_ULTRACLEANVETO_V3_v1.txt) is included to represent the isotropic gamma-ray background from various unresolved extragalactic sources. The analysis pipeline employed the Fermitools (v2.2.0) software within the Fermi Science Support Center’s Fermibottle Docker environment Wood et al. (2017).
@@ IV.2 Analysis Method
Our analysis followed the standard binned-likelihood method developed by the Fermi-LAT collaboration for likelihood calculations and model fitting Atwood et al. (2009). We incorporated sources from the 4FGL catalog up to beyond the defined region of interest into our model, freezing all their parameters to the catalog values. This was done to minimize bias from the potential presence of bright sources outside the targeted area and to account for the LAT’s poor point spread function (PSF) at low energies around 0.1 GeV Thorpe-Morgan et al. (2021); Manna and Desai (2024a). In the first phase of the analysis, the spectral parameters for all the free sources were determined. During the second phase, we then fixed the spectral parameters of all sources, except for the normalizations to their best-fit values. Then, we included a template for potential dark matter annihilation emission in the model which we modeled as a diffuse source with a spatial distribution proportional to the -factor calculated using CLUMPY and specified in Table LABEL:tab:TableII. The spectral component of the signal was given by the spectrum of annihilating dark matter for the specified channel utilizing the DMFitFunction Jeltema and Profumo (2008) within Fermitools similar to Thorpe-Morgan et al. (2021).
We used the gtapps tool for our analysis and employed the Maximum Likelihood Estimation (MLE) technique to identify the model parameters that best match the source’s spectrum and location. The gtlike tool performs a binned likelihood analysis on Fermi-LAT gamma-ray data. It achieves this by comparing a model of the gamma-ray sky with the actual observations and calculating the likelihood that the model explains the data. To identify gamma-ray sources and assess their significance, we employ the Test Statistic (TS) calculated using the gttsmap tool. This statistic, defined as Mattox et al. (1996)
compares the likelihoods of two scenarios: where is a model without the source (null hypothesis) and is a model including the source at a specific location (alternative hypothesis). Wilks’ theorem tells us that for high photon counts, the TS for the null hypothesis behaves similarly to a chi-squared distribution Wilks (1938). The detection significance is thereby nominally obtained by taking the square root of TS. However, D23 has shown using simulations, that for the null hypothesis the TS statistics has larger tails, which are not expected in the distribution. Therefore, the actual detection significance would be expected to be smaller than the square root of TS. Nevertheless, we shall quote the significance as the square root of TS for our results. We note that the same TS statistic is also used in neutrino and soft gamma-ray astrophysics to evaluate the significance of detections Pasumarti and Desai (2022); Manna and Desai (2024b).
@@ V Results
We conducted the analysis for the clusters in our sample using and annihilation channels. The TS values for all the 350 galaxy clusters can be found in Table LABEL:tab:TableII. Among all the clusters, SPT-CL J2021-5257 shows the maximum TS value of 9, corresponding to 3 emission using the dark matter template as shown in Figure 2 for the energy range between 1 GeV and 300 GeV. We also performed the analysis for SPT-CL J2021-5257 by extending the lower energy range to 100 MeV and found that TS value decreases to 7, corresponding to 2.6 significance. Other clusters, such as SPT-CL J2012-5649, SPT-CL J0124-5937, and SPT-CL J2300-5331 show TS values corresponding to significance, suggesting a weaker detection significance. The variation of TS values as a function of the dark matter mass for the aforementioned clusters with significance can be found in Figure 3 for annihilation channel . The best fit WIMP mass for SPT-CL J2021-5257 is found to be GeV, whereas the best-fit value of for this cluster assuming a annihilation channel. For SPT-CL J2012-5649, SPT-CLJ 0124-5937, and SPT-CL J2300-5331, the best-fit masses were found to be GeV, GeV, and GeV, respectively. The corresponding best-fit values of for the channel were , and , respectively. The best-fit dark matter mass for SPT-CL J2021-5257 for the annihilation channel is found to be GeV, whereas the best-fit annihilation cross-section is given by . Similarly, for SPT-CL J2012-5649, SPT-CLJ0124-5937, and SPT-CL J2300-5331 we found the best-fit mass as , , and GeV, respectively, whereas the best-fit values of are equal to , and , respectively for this annihilation channel. The variation of TS with mass for the annihilation channel for all the four clusters can be found in Fig. 4.
@@
As pointed out in D23, these values are in conflict with the upper limits obtained from null searches for dark matter annihilation from Milky Way dwarf spheroidal galaxies, whose 95% c.l. limits on the velocity-averaged annihilation cross-section are around Di Mauro and Winkler (2021). Therefore, although our significance does not cross the threshold, these enhanced TS values corresponding to (2-3) significance cannot be by-products of dark matter annihilation. These are either statistical fluctuations or caused by some astrophysical processes.
@@
Therefore, we calculate upper limits for the dark matter annihilation cross-section for these clusters. We show the 95% c.l. upper limits on the annihilation cross-section of WIMP dark matter mass for the annihilation channel in Figure 5 and for the channel in Figure 6. Our upper limits are mostly in agreement with other works, which have obtained limits on annihilation cross-section Di Mauro et al. (2023); Ackermann et al. (2010); Han et al. (2012); Thorpe-Morgan et al. (2021). Among all the clusters with significance , we found the most stringent limit for SPT-CL J2300-5331, viz. for GeV and annihilation channel and for GeV and annihilation channel. We note among the clusters with signifiance is SPT-CL J2012-5649 (TS = 6.2), which is spatially coincident with the merging cluster Abell 3667. We note that this cluster was detected with the highest significance () in our previous work on gamma-ray searches using a point source template Manna and Desai (2024a).
All other clusters have significance less then .
@@
These include the merging clusters in our sample such as Bullet Cluster (SPT-CL J0658-5556) and El Gordo Cluster (SPT-CL J0102-4915) with TS values of 1.5 and 1.9, respectively, corresponding to no significant excess. Since the last 100 clusters in our sample (sorted according to ) showed no significance we restricted our analysis to 350 clusters and did not go beyond that.
VI Conclusions
In this work, we searched for gamma-ray emission between 1 and 300 GeV from dark matter annihilation in galaxy clusters. We analyzed 350 clusters selected from the SPT-SZ 2500 square degree survey, which provides a mass-limited sample. The analysis utilized 15.7 years of data from the Fermi-LAT telescope and employed the DMFIT template. Notably, we identified a emission signal from the galaxy cluster SPT-CL J2021-5257. The best fit value for for SPT-CL J2021-5257 is found to be and best fit dark matter mass as GeV assuming annihilation channel and for annihilation channel we found the best fit value as and best fit dark matter mass as GeV. The TS map for this cluster can be found in Fig. 2. Although prima-facie, we found detection evidence from SPT-CL J2021-5257, the estimated annihilation cross-section is in conflict with upper limits from Milky way dwarf spheroidal galaxies. Therefore, we conclude that the enhanced signal seen for SPT-CL J2021-5257 cannot be due to dark matter annihilation. We obtained a marginal significance of around 2 for three other clusters: SPT-CL J2012-5649, SPT-CL J0124-5937 and SPT-CL J2300-5331. All the remaining clusters showed null results and the TS values were consistent with background.
The most stringent upper limits we obtain for annihilation to is given by SPT-CL J2300-5331 as for GeV and annihilation channel and for GeV and annihilation channel.
Acknowledgements.
We are grateful to the Fermi-LAT and CLUMPY teams for their invaluable contributions in making the data and analysis codes publicly available. Their expertise and willingness to answer our queries were instrumental in this research. This work was motivated after a stimulating talk by Mattia Di Mauro at IIT Hyderabad. We also acknowledge the National Supercomputing Mission (NSM) for providing essential computing resources on the ‘PARAM SEVA’ supercomputer at IIT Hyderabad. This initiative is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and the Department of Science and Technology (DST) of the Government of India. Finally, we thank the Smithsonian Astrophysical Observatory for developing SAOImageDS9, which played a valuable role in our analysis.References
- Cirelli (2012) M. Cirelli, Pramana 79, 1021 (2012), eprint 1202.1454.
- Donato (2014) F. Donato, Physics of the Dark Universe 4, 41 (2014).
- Chan (2021) M. H. Chan, Galaxies 9, 11 (2021).
- Desai (2018) S. Desai, Physics Letters B 778, 325 (2018), eprint 1708.06502.
- Zhao (1996) H. Zhao, MNRAS 278, 488 (1996), eprint astro-ph/9509122.
- Wilks (1938) S. S. Wilks, Annals Math. Statist. 9, 60 (1938).

### v2

@@ IV.1 Data selection
We also applied the filters, DATA_QUAL > 0 and LAT_CONFIG == 1, which were utilized to eliminate low-quality data. The rock angle was constrained to abs(rock angle) and |b| to mitigate particle contamination. To reduce atmospheric interference, a zenith angle cut of 90∘ was imposed on the events. We also removed all clusters within of the galactic plane.
The spatial binning of the data was performed at a resolution of pixels, which helps to reduce noise and improve the signal-to-noise ratio. Our background model incorporated all the gamma-ray sources from the fourth Fermi-LAT catalogue (4FGL-DR4) Ballet et al. (2023), encompassing both point-like and extended sources. To account for the pervasive, low-intensity gamma-ray emission from the Milky Way galaxy itself (diffuse emission), we employed the Galactic diffuse emission model (gll_iem_v07.fits). Additionally, an isotropic component (iso_P8R3_ULTRACLEANVETO_V3_v1.txt) is included to represent the isotropic gamma-ray background from various unresolved extragalactic sources. The analysis pipeline employed the Fermitools (v2.2.0) software within the Fermi Science Support Center’s Fermibottle Docker environment Wood et al. (2017).
@@ IV.2 Analysis Method
Our analysis followed the standard binned-likelihood method developed by the Fermi-LAT collaboration for likelihood calculations and model fitting Atwood et al. (2009). We incorporated sources from the 4FGL catalog up to beyond the defined region of interest into our model, freezing all their parameters to the catalog values. This was done to minimize bias from the potential presence of bright sources outside the targeted area and to account for the LAT’s poor point spread function (PSF) at low energies around 0.1 GeV Thorpe-Morgan et al. (2021); Manna and Desai (2024a). In the first phase of the analysis, the spectral parameters for all the free sources were determined. During the second phase, we then fixed the spectral parameters of all sources, except for the normalizations to their best-fit values. Then, we included a template for potential dark matter annihilation in the model, which we modeled as a diffuse source with a spatial distribution proportional to the -factor calculated using CLUMPY and specified in Table LABEL:tab:TableII. The spectral component of the signal was given by the spectrum of annihilating dark matter for the specified channel utilizing the DMFitFunction Jeltema and Profumo (2008) within Fermitools similar to Thorpe-Morgan et al. (2021).
We used the gtapps tool for our analysis and employed the Maximum Likelihood Estimation (MLE) technique to identify the model parameters that best match the source’s spectrum and location. The gtlike tool performs a binned likelihood analysis on Fermi-LAT gamma-ray data. It achieves this by comparing a model of the gamma-ray sky with actual observations and calculating the likelihood that the model explains the data. To identify gamma-ray sources and assess their significance, we employ the Test Statistic (TS) calculated using the gttsmap tool. This statistic, defined as Mattox et al. (1996)
compares the likelihoods of two scenarios: where is a model without the source (null hypothesis) and is a model including the source at a specific location (alternative hypothesis). Wilks’ theorem tells us that for high photon counts, the TS for the null hypothesis behaves similar to a chi-squared distribution Wilks (1938). The detection significance is thereby nominally obtained by taking the square root of TS. However, D23 has shown using simulations, that for the null hypothesis the TS statistics has larger tails, which are not expected in the distribution. Therefore, the actual detection significance would be expected to be smaller than the square root of TS. Nevertheless, we shall quote the significance as the square root of TS for our results. We note that the same TS statistic is also used in neutrino and soft gamma-ray astrophysics to evaluate the significance of detections Pasumarti and Desai (2022); Manna and Desai (2024b).
@@ V Results
We conducted the analysis for the clusters in our sample using the and annihilation channels in the energy range between 1-300 GeV. The TS values for all the 350 galaxy clusters (for the annihilation channel) can be found in Table LABEL:tab:TableII. Among all the clusters, SPT-CL J2021-5257 shows the maximum TS value of 9.2, corresponding to 3.03 significance. The corresponding TS map for this cluster is shown in Figure 2. We also performed the analysis for SPT-CL J2021-5257 by extending the lower energy range to 100 MeV, and found that the TS value decreases to 7 corresponding to 2.6 significance. Other clusters, such as SPT-CL J2012-5649, SPT-CL J0124-5937, and SPT-CL J2300-5331 show TS values corresponding to significance, suggesting a weaker detection significance. The variation of TS values as a function of the dark matter mass for the aforementioned clusters with significance can be found in Figure 3 for annihilation channel. The best fit WIMP mass for SPT-CL J2021-5257 is found to be GeV, whereas the best-fit value of for this cluster assuming a annihilation channel. For SPT-CL J2012-5649, SPT-CLJ 0124-5937, and SPT-CL J2300-5331, the best-fit masses were found to be GeV, GeV, and GeV, respectively. The corresponding best-fit values of for the channel were , and , respectively. The best-fit dark matter mass for SPT-CL J2021-5257 for the annihilation channel is found to be GeV, whereas the best-fit annihilation cross-section is given by . Similarly, for SPT-CL J2012-5649, SPT-CLJ0124-5937, and SPT-CL J2300-5331 we found the best-fit mass as , , and GeV, respectively, whereas the best-fit values of are equal to , and , respectively for this annihilation channel. The variation of TS value with the WIMP mass for the annihilation channel for all the four clusters can be found in Fig. 4. We also show how TS varies with other subhalo models and uncertainties on -factors in Appendix A, B, and C.
@@
As pointed out in D23, these values are in conflict with the upper limits obtained from null searches for dark matter annihilation from Milky Way dwarf spheroidal galaxies, whose 95% c.l. limits on the velocity-averaged annihilation cross-section are around Di Mauro and Winkler (2021). Therefore, although our significance does not cross the threshold, these enhanced TS values corresponding to (2-3) significance cannot be by-products of dark matter annihilation. Furthermore, after considering the look elsewhere effect, given that we have analyzed 350 galaxy clusters Gross and Vitells (2010), the global -value for the cluster with the highest TS value of around 9.2, is equal to 0.47, which is consistent with the expected background. However, if we ignore the look elsewhere effect, and in case the significance increases with additional exposure, then one explanation is that this signal could be due to some astrophysical process resulting from cosmic interactions with gas and photons in the intracluster medium, as pointed out in D23.
@@
Therefore, we calculate upper limits for the dark matter annihilation cross-section for these clusters. For the aforementioned four clusters, we show the 95% c.l. upper limits on the annihilation cross-section as a function of WIMP mass for the annihilation channel in Figure 5 and for the channel in Figure 6. Our upper limits are mostly in agreement with other works, which have obtained upper limits on the annihilation cross-section Di Mauro et al. (2023); Ackermann et al. (2010); Han et al. (2012); Thorpe-Morgan et al. (2021). A compilation of upper limits for all the remaining galaxy clusters analyzed in this work can be found in Table LABEL:tab:TableIII, as discussed in Sect. VI. Among all the clusters with significance , we found the most stringent limit for SPT-CL J2300-5331, viz. for GeV considering the annihilation channel, and for GeV and considering the annihilation channel. We note that one of the clusters with significance is SPT-CL J2012-5649 (TS = 6.4), which is spatially coincident with the merging cluster Abell 3667. This cluster was detected with the highest significance () in our previous work on gamma-ray searches using a point source template Manna and Desai (2024a).
@@
All other clusters have significance less then . These include the merging clusters in our sample such as Bullet Cluster (SPT-CL J0658-5556) and El Gordo Cluster (SPT-CL J0102-4915) with TS values of 1.5 and 1.9, respectively, corresponding to no significant excess. Since the last 100 clusters in our sample (sorted according to ) showed no significance we restricted our analysis to 350 clusters and did not go beyond that.